Given that 79=2 (mod x) where 0 < x <100. Find the set of values for which the statement is true

THANK YOU

Answer by ikleyn(52781)

.
Given that 79=2 (mod x) where 0 < x <100. Find the set of values for which the statement is true
~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~

The fact that 79 = 2 (mod x) means that 79-2 = 0 (mod x),  which is the same as 77 = 0 (mod x).

In turn, it means that 77 is divisible by x.
But 77 has only two divisors: 7 and 11. 

Yes, Seven and Eleven.

Answer.  x is 7 or 11.